OpenStudy (divinexxchocolatexx):

Please help me?***WILL MEDAL AND FAN*** Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 −2x + y2 − 6y = 26

OpenStudy (sshayer):

either complete the squares and write in the form (x-h)^2+(y-k)^2=r^2 center is (h,k) and radius =r second method compare with \[x^2+y^2+2gx+2fy+c=0\] center is (-g,-f) and radius \[r=\sqrt{g^2+f^2-c}\]

satellite73 (satellite73):

you know how to complete the square?

OpenStudy (divinexxchocolatexx):

@satellite73 nope

OpenStudy (divinexxchocolatexx):

@sshayer i dont understand anything you said because of the way you worded it

OpenStudy (mathmale):

Have you read any explanations (in your book or online) of how to find the center and vertex of a circle? about how to "complete the square?" If not, now would be a good time to look up those two topics.

OpenStudy (divinexxchocolatexx):

@mathmale ok thanks

OpenStudy (divinexxchocolatexx):

@mathmale i sarched it up but it doesnt make any sense

OpenStudy (divinexxchocolatexx):

*searched

Nnesha (nnesha):

chocolates! first i'm going to put parentheses to separate x's and y's terms \[\rm (x^2 −2x )+( y^2 − 6y )= 26\] now we have to complete the square for each parentheses

Nnesha (nnesha):

take half of the middle term(coefficient of x term) square it and then add this value to both side here is an example \[(z^2-6z)= 4\] take half of middle term (which is 6 in this example ) 6/2= 3 now square this value (3)^2 = 9 add the resultto the both sides \[(z^2-6z\color{Red}{+9})= 4\color{red}{+9}\]

Nnesha (nnesha):

you can factor that quadratic equation at left side but in short form the factor of the trinomial that you get after completing the square would always be \[\rm (x-\frac{b}{2})^2\]

OpenStudy (divinexxchocolatexx):

wait.. I dont under stand the (z^2- 6z part could you explain it again

Nnesha (nnesha):

sure. thanks for asking

OpenStudy (divinexxchocolatexx):

btw thank you sooo much for helping

Nnesha (nnesha):

\[\large\rm Ax^2+Bx+C=0\] the quadratic equation where `b` is the middle term ( the coefficient of x term) first we need to see if the leading coefficient is one (if not we need to solve the equation to get rid of the leading coefficient ) to complete the square you have to take half of `b` ( coefficient of x term )

Nnesha (nnesha):

\[\large\rm x^2+Bx+C=0\] for example this is my equation 1st) leading coefficient is one ! 2nd) move the constant term to the right side \[\large\rm \color{blue}{x^2+Bx} =-C\] complete the square of blue part 3rd) take half of middle term ( b) and then square it ( add the result both side of the equation ) \[\rm x^2+Bx +\color{Red}{(\frac{b}{2})} =-C \color{Red}{+(\frac{b}{2})^2}\] (in simple words { divide b by 2, square it , add this value both sides )

OpenStudy (divinexxchocolatexx):

ohh ok

Nnesha (nnesha):

there is a typo

OpenStudy (divinexxchocolatexx):

where

Nnesha (nnesha):

\[\rm x^2+Bx +\color{Red}{(\frac{b}{2})^2} =-C \color{Red}{+(\frac{b}{2})^2}\] **

OpenStudy (divinexxchocolatexx):

oh

OpenStudy (divinexxchocolatexx):

Thank you!

Nnesha (nnesha):

\[\rm \color{red}{x^2+Bx +}\color{Red}{(\frac{b}{2})^2} =-C \color{black}{+(\frac{b}{2})^2}\] and then factor the left side which will always be \[\rm (x+\frac{b}{2})^2 \] (x+half of middle term )^2

Nnesha (nnesha):

np. can you complete the square of `x^2-2x` ?

OpenStudy (divinexxchocolatexx):

I think so

OpenStudy (divinexxchocolatexx):

2/2=1 1*1=1 so just add the 1 right?

Nnesha (nnesha):

okay go ahead give it a try!

Nnesha (nnesha):

yes that's correct. btw `x^2-2x` b=-2 take half -2/2 = -1 and when you take square of negative number you will always get the positive answer. nice

OpenStudy (divinexxchocolatexx):

You are a life saver!

Nnesha (nnesha):

what would be the next step? write the the equation form

Nnesha (nnesha):

in**

OpenStudy (divinexxchocolatexx):

Would i do the same thing for the other parentheses?

OpenStudy (divinexxchocolatexx):

since i have two numbers to add on which do i add on the 26

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @DivinexxChocolatexx Would i do the same thing for the other parentheses? \(\color{blue}{\text{End of Quote}}\) not yet we still have to work on first parentheses add (b/2)^2 both sides \[(\color{Red}{x^2 -2x+1}) +(y^2-6y)=26+(-1)^2\]what are the factors of this perfect square trinomial ?

OpenStudy (divinexxchocolatexx):

tbh I don't know

OpenStudy (divinexxchocolatexx):

wait is it 2 and 6?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\rm \color{red}{x^2+Bx +}\color{Red}{(\frac{b}{2})^2} =-C \color{black}{+(\frac{b}{2})^2}\] and then factor the left side which will always be \[\rm (x+\frac{b}{2})^2 \] (x+half of middle term )^2 \(\color{blue}{\text{End of Quote}}\)

Nnesha (nnesha):

x^2-2x+1 when the leading coefficient is one then find two numbers when you multiply them you should get product of `A times C` and the sum of same numbers should be equal to `b` ( the middle term )

Nnesha (nnesha):

Ax^2+Bx+C=0 quadratic equation where a=leading coefficient b= coefficient of x term( middle term) c=constant

OpenStudy (divinexxchocolatexx):

2 and 26?

Nnesha (nnesha):

first of all x^2-2x+1 list a , b ,c values

OpenStudy (divinexxchocolatexx):

a value is 2 and b value is 6 and c value is 26

Nnesha (nnesha):

hmm `x^2-2x+1` i don't see any `26`

OpenStudy (divinexxchocolatexx):

ohhhh

OpenStudy (divinexxchocolatexx):

so c value is 1

Nnesha (nnesha):

\[\rm (x^2 −2x )+( y^2 − 6y )= 26\] we are completing the square of first parentheses \[\rm (x^2 −2x +\color{Red}{(\frac{-2}{2})^2 })+( y^2 − 6y )= 26\color{Red}{+(\frac{-2}{2})^2}\] add (b/2)^2 both sides we get \[\rm (x^2 −2x +\color{Red}{ 1 })+( y^2 − 6y )= 26\color{Red}{+1}\]

Nnesha (nnesha):

add 26+1 =27 \[\rm \color{blue}{(x^2 −2x +\color{Red}{ 1 })}+( y^2 − 6y )= 27 \] now we have to factor this trinomial(quadratic equation ) so at this point we are only dealing with the blue part (quadratic equation

Nnesha (nnesha):

yes c=1 a=1 b=-2 multiply A times C = 1 times 1 what two numbers would you multiply to get 1( the product of AC) and when you add these same numbers you should get (-2) (the value of b)

OpenStudy (divinexxchocolatexx):

omg that makes so much sense

OpenStudy (divinexxchocolatexx):

this is a totally random question but what time is it where you are

Nnesha (nnesha):

we don't use clocks where i live c;

OpenStudy (divinexxchocolatexx):

lmao so what you use

Nnesha (nnesha):

at this point you have to learn how to factor quadratic equation but for this question when we complete the square, the factor of trinomial would always be (x+half of b)^2

OpenStudy (divinexxchocolatexx):

oh i know how to do that

Nnesha (nnesha):

\[\rm (x^2 −2x )+( y^2 − 6y )= 26\] we are completing the square of first parentheses \[\rm (x^2 −2x +\color{Red}{(\frac{-2}{2})^2 })+( y^2 − 6y )= 26\color{Red}{+(\frac{-2}{2})^2}\] add (b/2)^2 both sides we get \[\rm (\color{Red}{x^2 −2x +\color{Red}{ 1} })+( y^2 − 6y )= 27\] \[\rm (\color{Red}{(x+\frac{b}{2})^2 })+( y^2 − 6y )= 27\] replace b with its value that's it and then simplify

OpenStudy (divinexxchocolatexx):

to factor it?

Nnesha (nnesha):

well first just replace `b` with its value

OpenStudy (divinexxchocolatexx):

oh i thought i would have to times the first set of parentheses by itself

OpenStudy (divinexxchocolatexx):

Thank you for helping me but i need to go to sleep it's 1:01 in the morning

Nnesha (nnesha):

no no. that's kinda short cut let me explain it first `x^2-2x+1` factor this equation the product of leading coefficient and constant is 1 middle term is -2 two numberes are -1 and -1 -1 times -1 = 1 ( the product of AC) -1-1=-2 ( sum is equal to the middle term the factor are (x-1)(x-1) (x + 1st number )(x+2nd ) number

OpenStudy (divinexxchocolatexx):

good night :)

Nnesha (nnesha):

the short cut i was talking about earlier is (x+(b/d))^2 \[\rm (x^2 −2x )+( y^2 − 6y )= 26\] we are completing the square of first parentheses \[\rm (x^2 −2x +\color{Red}{(\frac{-2}{2})^2 })+( y^2 − 6y )= 26\color{Red}{+(\frac{-2}{2})^2}\] add (b/2)^2 both sides we get \[\rm (\color{Red}{x^2 −2x +\color{Red}{ 1} })+( y^2 − 6y )= 27\] \[\rm (\color{Red}{(x+\frac{b}{2})^2 })+( y^2 − 6y )= 27\] replace b with its value that's it and then simplify b =-2 plugin \[\rm (\color{Red}{(x+\frac{-2}{2})^2 })+( y^2 − 6y )= 27\] simplify -2/2 = -1 \[\rm (\color{Red}{x-1 })^2+( y^2 − 6y )= 27\] same answer you can say (x+b/d)^2 is the formula to factor trinomial when we complete the square

Nnesha (nnesha):

follow same steps to complete the square for 2nd parentheses! practice! and practice on your ow the lag is real _-- alright i'm tired too good night !